OpenStudy (anonymous):

A panel of judges must consist of four women and three men. A list of potential judges includes six women and five men. How many different panels could be created from this list?

OpenStudy (anonymous):

From those 6 women you choose 4 and from 5 men choose 3. You should use this\[\left(\begin{matrix}6 \\ 4\end{matrix}\right) \left(\begin{matrix}5 \\ 3\end{matrix}\right)\]

OpenStudy (anonymous):